Stuart Manufacturing Company established the following standard price and cost data.

Sales price $ 8.60 per unit
Variable manufacturing cost $ 3.80 per unit
Fixed manufacturing cost $ 2,300 total
Fixed selling and administrative cost $ 800 total

Stuart planned to produce and sell 2,100 units. Actual production and sales amounted to 2,300 units.

Required

  1. Determine the sales and variable cost volume variances.

  2. Classify the variances as favorable (F) or unfavorable (U).

  1. Determine the amount of fixed cost that will appear in the flexible budget.

  2. Determine the fixed cost per unit based on planned activity and the fixed cost per unit based on actual activity.
Answer #1

Ans:

A). Sales variance= Actual sales unit - budgeted sales unit)* Contribution per unit {sales price- variable manufacturing cost}

= 2300-2100)* (8.60-3.80)

=200*4.80 =960 favorable

Variable cost variance= Budgeted units produced - actual units produced }* variable cost per unit

=(2100-2300)* 3.8

=760 unfavorable

B). Variances are classified as favorable or unfavorable in A part only

Sales variance is said to be favorable variance because here the number of units produced was more than the budgeted units planned to produce

Variable cost variance is said to be unfavorable variance because here actual units are more than standard

C). Amount of fixed cost that will appear in flexible budgeted

For Fixed manufacturing cost= Fixed cost/ actual cost

=2300/2300 = $1.00

For Fixed selling and administrative cost= fixed selling/ actual cost

800/2300 =$0.34

D). Fixed per unit on planned activity= total cost/2100 units

=2300+800/2100

=3100/2100=1.47

Fixed per unit on budgeted/ actual activity = 3100/2300

=$1.34